OpenStudy (anonymous):

The isotope Cobalt-60 has a nuclear mass of 59.933820 u Calculate the Mass Defect of Cobalt-60 using the following information. Mass of Proton: 1.007825 u Mass of Neutron: 1.008665 u 1 u = 931.5 MeV

OpenStudy (anonymous):

cobalt has 27 protons and 33 neutrons in its nucleus so M=27*1.007825u+33*1.008665u =27.211u+33.285u =60.496u But the measured mass of co-60 is given as 59.933820 u \[DeltaM=M-measured mass\] \[DeltaM=60.496u-59.933820 u\] MASS DEFECT =0.56218u = 0.56218*931.5 Mev MASS DEFECT =523.670 Mev